Your banker has analyzed your company account and has suggested that her bank has a cash management package for you. She suggests that with a concentration banking system, your float can be reduced by four days on average. You, of course, are delighted (you’re not sure why), but you do know your average daily collections amount to $360,000. Your opportunity cost of funds is 8 percent. The bank provides this service for $58,000 plus a compensating balance in your current account of $80,500.

1. is this package worth it?

2. by how much? (annual saving)
Answer #1

Step 1

We will first find the cost of funds saved by reducing the float days

=Average daily collection * Float days reduced*cost of funds

=$360,000*4*8%

=$115,200

Step 2

We will find the cost of package

=Bank Service fee+cost of funds on compensating balance

=$58,000 +($80,500*8%)

=$64,440

a)YES, the package should be accepted as the cost saved by package(step 1) is higher than cost of package.(step 2)

b) Annual savings = step 1 -step 2

=$115,200-$64,440

=$50,760

Cost will be saved by $50,760

*Compensating balance is the amount required to maintain interest free at the bank.
Answer #2

SOLUTION :


Collection in one day = $360000 .

Collection in 4 days = $360000 * 3 = $961440000 .

Collection period is reduced by 4 days means cash of $1440000 will not be blocked in the system and will be available to the company.


So, 


Saving of interest on this cash-increase in cash availability throughout the year

= 1440000 * 0.08

= 115200 ($).


Cost of getting this system 

= Fees of the bank + loss of interest on compensating balance

= 58000 + 80500*0.08

= 64440 ($)


1.


The package offered by the bank is  worth taking since savings are much more than the cost. (ANSWER).


2.


The extent of gains (savings) annually :

= Savings - Cost 

= 115200 - 64440 

= 50760 ($)  (ANSWER).

answered by: Tulsiram Garg